OpenStudy (anonymous):

Estimate Square root of 45 to the nearest tenth.

OpenStudy (anonymous):

 \[\sqrt{5}\] has to be greater than 2 because 2*2=4 but less than 2.3 because 2.3* 2.3=5.29 so square root three is approximatley 2.2 and 2.2*3=6.6